OpenStudy (anonymous):

When solving the system of equations by graphing, what is the solution of 3x + 2y = 2 and 2x – y = 6?

OpenStudy (lncognlto):

3x + 2y = 2 ...(1) 2x - y = 6 ... (2) (2) x 2: 4x - 2y = 12 + (1): 3x + 2y = 2 therefore 7x = 14 therefore x = 2 therefore y = -2 (2,-2) is the solution of the equations